Can someone help me with the following problem: Solve the following systems of equatons algebraically. Be certain you have solved both variables in teh frist sytem of equations and all 3 variables in teh second system. 1, 5x+y = 5
3x-5y = -63
2. a+b+3c= - 8
2b-c=15
3a+2c=7.
This is the answer I got am I on the right track? 1. x= 4 and y = -1 2. a=77, b -8 and c= - 30. If I am wrong can you show me how you solved it? Thank you!!!!

Answer by stanbon(48558)

Can someone help me with the following problem: Solve the following systems of equatons algebraically. Be certain you have solved both variables in teh frist sytem of equations and all 3 variables in teh second system.
1,
5x+y = 5
3x-5y = -63
---
Your solution is incorrect.
5x+y = 5
3x-5y = -63
--
Solving 1st equation for "y" you get y = -5x+5
Substituting that into the 2nd equation you can solve for "x":
3x -5(-5x+5) = -63
3x +25x-25 = -63
28x = 38
x = 19/14
---
Substitute that into y = -5x+5 = -5(19/14)+5 = -95/14 + 70/14 = -25/14
Solution: x = (19/14) ; y = (-25/14)
------------------------
2.
a + b+3c =-8
0a+2b-c =15
3a+0b+2c =7
------------------------
From the 2nd equation b = (1/2)c + (15/2)
From the 3rd equation a = (-2/3)c + (7/3)
--------------------------------------------
Substitute those into the 1st equation and solve for "c":
(-2/3)c + (7/3) + (1/2)c + (15/2) + 3c = -8
Multiply thru by 6 to get:
-4c + 14 + 3c + 45 +18c = -48
17c = -107
c = -107/17
-------------------
Substitute that into b=(1/2)c+(15/2) to get b = 74/17
----
Substitute the value for c into a=(-2/3)c + (7/3) to get a = 111/17
======================================================================
Cheers,
Stan H.




This is the answer I got am I on the right track?
1. x= 4 and y = -1
2. a=77, b -8 and c= - 30.
If I am wrong can you show me how you solved it?